# Parallel GPU-Enabled Algorithms for SpGEMM on Arbitrary Semirings with Hybrid Communication

### v1

@@ Abstract.
In this paper, we implement a GPU-based distributed-memory SpGEMM routine on top of the CombBLAS library. Our implementation achieves a speedup of over 2×2$\times$2 × compared to the CPU-only CombBLAS implementation and up to 3×3$\times$3 × compared to PETSc for large input matrices.
@@ 1. Introduction
Given matrix multiplication underlies a significant part of HPC applications, considerable efforts have been made in the last three decades to optimize dense matrix multiplication on sequential and parallel machines (Van De Geijn and Watts, 1997; Goto and Geijn, 2008; Irony et al., 2004). However, many scientific and engineering disciplines rely on Sparse General Matrix Multiply (SpGEMM), where two sparse input matrices are multiplied to produce a sparse output matrix. For example, each graph can be represented as a sparse adjacency matrix, where the entries denote the edges between the nodes (Zwick, 2002). Problems ranging from quantum chemistry to genomics to graph algorithms can be formulated as sparse graphs and thus as adjacency matrices (Matam et al., 2012; Guidi et al., 2021; Gleinig, 2022) and processed with SpGEMM.
The expressive capabilities of SpGEMM are limited by the fixed, precise nature of matrix multiplication (the multiplication, the addition and the type of the entries or nonzeros). The semiring abstraction enables user-defined redefinition of entry type, multiplication and addition operator (Glazek, 2002). Using the semiring abstraction, numerous graph operations such as breadth first search or $k$-edge shortest path can be reduced to a sparse matrix multiplication over a semiring (Burkhardt, 2021).
Graph sizes have increased rapidly in recent years, for example in genomics with new sequencing technologies (Chen et al., 2021). The adjacency matrices of these graphs can exceed single-node memory capacities, so distributed memory parallelism is required to avoid I/O performance degradation. On the one hand, CombBLAS provides a set of linear algebra primitives with semiring support, parallelized over distributed memory (Buluç and Gilbert, 2011). However, CombBLAS does not currently provide GPU parallelism in its routines, with the exception of one application implemented with the library, HipMCL (Selvitopi et al., 2020), which does not require semiring support. On the other hand, GALATIC is a library that supports SpGEMM on custom semirings on a single GPU (Lett, 2021).
@@ 2.1. Distributed SpGEMM
SpGEMM is a widely-used kernel for multiplying sparse matrices, often found in graph applications where the adjacency matrices consist mainly of zeros. Dense matrix multiplication is inefficient for sparse matrices, so more memory-efficient, specialized storage formats such as CSR, CSC or DCSC are used for sparse matrices (Willcock and Lumsdaine, 2006).
In graph applications, sparse matrices tend to grow larger than the memory capacity of a single node. The memory requirement is $\mathcal{O}(n^{2})$ with respect to the vertices. Given this order of magnitude, even high-end computing nodes reach their memory capacity and distributed-memory parallelism is required. In distributed SpGEMM algorithms, multiple processes handle parts of the multiplication and communicate the results together. In general, local computation within a node is parallelized via OpenMP (Dagum and Menon, 1998) and communication between nodes is handled via MPI (Walker and Dongarra, 1996).
On distributed systems, we execute SpGEMM via the Sparse Scalable Universal Matrix Multiplication Algorithm (Sparse SUMMA) (Buluç and Gilbert, 2011; Van De Geijn and Watts, 1997). This algorithm divides the processes into a 2D process grid. The CombBLAS implementation of Sparse SUMMA requires a square number of processes, but the algorithm itself does not. The multiplication can be regarded as a dot product of subranges of the input matrices $\mathbf{A}$ and $\mathbf{B}$, corresponding to a subrange of the output matrix $\mathbf{C}$. The multiplication is performed by multiplying each process simultaneously over its row and its column, each process being the broadcaster and the receiver of a sub-range of the matrix. The broadcast replicates the relevant local sub-matrix over the entire row or column of the process grid.
@@
CombBLAS uses a 2.5D algorithm to perform the multiplication. The .5.5.5.5 indicates that each process divides its matrix, resulting in two rounds of multiplication. The $\mathbf{A}$ matrix is halved column-wise, and the $\mathbf{B}$ matrix is halved row-wise. The actual multiplication is performed by local routines, which reduces the size of the matrices and allows larger matrices to be stored in less memory. The results are then merged into a product matrix.
@@ 2.2. Semiring Abstraction
In a nutshell, the semiring abstraction enables the user to redefine the type of matrix entries, multiplication and addition operator. By definition, a semiring is an algebraic structure that does not require additive inverse (Goodman, 1999). This definition provides a minimal set of axioms with which matrix multiplication can still be performed, since it requires only commutative addition, the existence of zero and one, and associative operations. It only requires commutative addition, the existence of zero and one, and associative operations and allows the user to redefine addition and multiplication operations.
The semiring abstraction makes it possible to, for example, formulate graph algorithms as matrix multiplication. In the shortest path problem, we recast the multiplication as addition and addition as the $\min$. Considering the adjacency matrix of any graph $W$, $W^{k}$ is the shortest path between vertices using $k$ vertices, and $W^{n}$ is the distance matrix (Zwick, 2002).
@@ 2.3. CombBLAS and GALATIC
CombBLAS is the primary library used to implement the algorithms in this paper (Buluç and Gilbert, 2011). It provides a high-performance implementation of various sparse linear algebra kernels with support for user-defined semirings. CombBLAS uses the Compressed Sparse Column (CSC) or Doubly Compressed Sparse Column (DCSC) format to store matrices. However, CombBLAS does not currently support GPU computation.
In contrast, GALATIC (Lett, 2021) provides local SpGEMM on GPU and support for semiring abstraction but does not implement distributed memory parallelism. By extending the existing AC SpGEMM library (Winter et al., 2019), GALATIC achieves a significant speed-up compared to the local SpGEMM algorithms of CombBLAS. GALATIC uses a four-stage process for its operation. The process begins by dividing the non-zeros of $\mathbf{A}$ evenly among the blocks of a GPU. Then, the products are computed, expanded, sorted, and combined on the individual GPU threads. In the third step, parts of the product matrix are combined. The rows are then merged across the parts before the matrix is returned. This division of work makes for an extremely efficient SpGEMM, resulting in an algorithm that is faster than cuSPARSE for highly sparse matrices (Winter et al., 2019).
@@ 3. Related Work
Buluç et al. (Buluç and Gilbert, 2012) developed and compared several distributed SpGEMM algorithms, including the Sparse SUMMA algorithm used in this work. Azad et al. (Azad et al., 2016) presented a communication avoiding distributed SpGEMM algorithm that uses a 3D process grid and has a lower communication cost compared to Sparse SUMMA. Hong et al. (Hong and Buluc, 2024) presented a sparsity-aware distributed SpGEMM algorithm that reduces the communication cost by communicating only the non-zeros involved in the computation. These works do not consider the use of GPUs or explore hybrid communication.
Brock et al. (Brock et al., 2024) present a number of RDMA-based algorithms for distributed SpGEMM on GPUs, but do not investigate arbitrary semirings or hybrid communication. A large body of existing work (Dalton et al., 2015; Nagasaka et al., 2017; Winter et al., 2019) focuses on optimizing SpGEMM on single GPUs. Other works (Park et al., 2023; Liu and Vinter, 2015) present strategies for using CPUs and GPUs on the same node in lockstep to perform SpGEMM. However, these works do not target distributed or multi-GPU scenarios.
@@ 4.3. GALATIC SpGEMM
[…]
@@ 5. Results and Discussion
The experiments were performed on the GPU nodes of the Perlmutter Supercomputer at NERSC (NERSC, [n. d.]) using Cray-MPICH and CUDA Toolkit 12.4. Each GPU node is equipped with an AMD EPYC 7763 CPU, four NVIDIA A100 GPUs and four Cassini-1 Network Interface Cards (NICs). The GPUs on the same node are fully connected via 4 NVLink 3.0. The different nodes are connected via a 3-hop Dragonfly topology with an HPE Slingshot 11 interconnect.
The GPUs and CPUs are warmed up with an algorithm run that is not included in the statistics. I/O is excluded from the specified runtimes. The values for the first run can be found in Table 1. In GALATIC we used the default settings, except for MaxChunksGeneralizedMerge, which was set to 512 instead of 256 to handle larger matrices. The tests were performed with different matrices from the SuiteSparse Matrix Collection (Davis, 2019), where each operation was a multiplication of a matrix by itself over the semiring of float types. Relevant statistics can be found in Table 2.
@@ 5.1. Distributed GPU-Enabled SpGEMM
For each of the four matrices, we performed three experiments, each of which was repeated four times across four different process grid sizes. For external comparison, we used the PETSc library(Balay et al., 2024). PETSc offers 1D SpGEMM with row-wise partitioning and GPU acceleration. For comparison, we tested each matrix using CombBLAS CPU, CombBLAS GPU (this work), and PETSc GPU and performed $\mathbf{A}^{2}$ for each matrix. The tests ran on four nodes, with one GPU per MPI process. Our GPU routine used only CUDA-aware MPI broadcasts during these tests. The PETSc results in Figure 6 and the one process experiment in Figure 5 were excluded due to out of memory issues with large matrices. Our GPU-enabled CombBLAS version showed a significant speedup, although the magnitude varied depending on the number of processes. The highest speedup was 2×2$\times$2 × on the Long_Coup_dt0 matrix. Overall, PETSc was faster with smaller matrices, but is slower or unable to run on larger matrices.
@@ 6. Conclusions
In this paper, we presented a scalable GPU-enabled, distributed SpGEMM routine that support an arbitrary semiring. To achieve this goal, we integrated the GALATIC library into CombBLAS to effectively handle local SpGEMMs on the GPU. In addition, we modified the communication routines to operate on matrices resident in GPU memory. In this way, we were able to achieve a 1.5-2.5$\times$ speedup compared to the native CPU version of CombBLAS and up to 3$\times$ speedup compared to the GPU SpGEMM routine in PETSc.
Finally, we showed that inter-process communication can be faster when copied to the CPU and CPU-CPU communication is used instead of direct GPU-GPU communication via CUDA-aware MPI. Based on this insight, we implemented a hybrid communication scheme based on message sizes. This resulted in a speedup of the communication runtime between 2$\times$ and 4$\times$.

### v2

@@ Abstract.
In this paper, we implement a GPU-based distributed-memory SpGEMM routine on top of the CombBLAS library. Our implementation achieves a speedup of over $2\times$ compared to the CPU-only CombBLAS implementation and up to $3\times$ compared to PETSc for large input matrices.
@@ 1. Introduction
Given matrix multiplication underlies a significant part of HPC applications, considerable efforts have been made in the last three decades to optimize dense matrix multiplication on sequential and parallel machines (van1997summa; goto2008anatomy; irony2004communication). However, many scientific and engineering disciplines rely on Sparse General Matrix Multiply (SpGEMM), where two sparse input matrices are multiplied to produce a sparse output matrix. For example, each graph can be represented as a sparse adjacency matrix, where the entries denote the edges between the nodes (10.1145/567112.567114). Problems ranging from quantum chemistry to genomics to graph algorithms can be formulated as sparse graphs and thus as adjacency matrices (6507483; guidi2021parallel; gleinig2022optimizing) and processed with SpGEMM.
The expressive capabilities of SpGEMM are limited by the fixed, precise nature of matrix multiplication (the multiplication, the addition and the type of the entries or nonzeros). The semiring abstraction enables user-defined redefinition of entry type, multiplication and addition operator (glazek2002guide). Using the semiring abstraction, numerous graph operations such as breadth first search or $k$-edge shortest path can be reduced to a sparse matrix multiplication over a semiring (10.1145/3446216).
Graph sizes have increased rapidly in recent years, for example in genomics with new sequencing technologies (10.1016/j.gpb.2021.08.001). The adjacency matrices of these graphs can exceed single-node memory capacities, so distributed memory parallelism is required to avoid I/O performance degradation. On the one hand, CombBLAS provides a set of linear algebra primitives with semiring support, parallelized over distributed memory (bulucc2011combinatorial). However, CombBLAS does not currently provide GPU parallelism in its routines, with the exception of one application implemented with the library, HipMCL (selvitopi2020optimizing), which does not require semiring support. On the other hand, GALATIC is a library that supports SpGEMM on custom semirings on a single GPU (Lett_2021).
@@ 2.1. Distributed SpGEMM
SpGEMM is a widely-used kernel for multiplying sparse matrices, often found in graph applications where the adjacency matrices consist mainly of zeros. Dense matrix multiplication is inefficient for sparse matrices, so more memory-efficient, specialized storage formats such as CSR, CSC or DCSC are used for sparse matrices (dcsr).
In graph applications, sparse matrices tend to grow larger than the memory capacity of a single node. The memory requirement is $\mathcal{O}(n^{2})$ with respect to the vertices. Given this order of magnitude, even high-end computing nodes reach their memory capacity and distributed-memory parallelism is required. In distributed SpGEMM algorithms, multiple processes handle parts of the multiplication and communicate the results together. In general, local computation within a node is parallelized via OpenMP (openmp) and communication between nodes is handled via MPI (walker1996mpi).
On distributed systems, we execute SpGEMM via the Sparse Scalable Universal Matrix Multiplication Algorithm (Sparse SUMMA) (SparseSumma; van1997summa). This algorithm divides the processes into a 2D process grid. The CombBLAS implementation of Sparse SUMMA requires a square number of processes, but the algorithm itself does not. The multiplication can be regarded as a dot product of subranges of the input matrices $\mathbf{A}$ and $\mathbf{B}$, corresponding to a subrange of the output matrix $\mathbf{C}$. The multiplication is performed by multiplying each process simultaneously over its row and its column, each process being the broadcaster and the receiver of a sub-range of the matrix. The broadcast replicates the relevant local sub-matrix over the entire row or column of the process grid.
@@
CombBLAS uses a 2.5D algorithm to perform the multiplication. The .5.5 indicates that each process divides its matrix, resulting in two rounds of multiplication. The $\mathbf{A}$ matrix is halved column-wise, and the $\mathbf{B}$ matrix is halved row-wise. The actual multiplication is performed by local routines, which reduces the size of the matrices and allows larger matrices to be stored in less memory. The results are then merged into a product matrix.
@@ 2.2. Semiring Abstraction
In a nutshell, the semiring abstraction enables the user to redefine the type of matrix entries, multiplication and addition operator. By definition, a semiring is an algebraic structure that does not require additive inverse (goodman1999semiring). This definition provides a minimal set of axioms with which matrix multiplication can still be performed, since it requires only commutative addition, the existence of zero and one, and associative operations. It only requires commutative addition, the existence of zero and one, and associative operations and allows the user to redefine addition and multiplication operations.
The semiring abstraction makes it possible to, for example, formulate graph algorithms as matrix multiplication. In the shortest path problem, we recast the multiplication as addition and addition as the $\min$. Considering the adjacency matrix of any graph $W$, $W^{k}$ is the shortest path between vertices using $k$ vertices, and $W^{n}$ is the distance matrix (10.1145/567112.567114).
@@ 2.3. CombBLAS and GALATIC
CombBLAS is the primary library used to implement the algorithms in this paper (bulucc2011combinatorial). It provides a high-performance implementation of various sparse linear algebra kernels with support for user-defined semirings. CombBLAS uses the Compressed Sparse Column (CSC) or Doubly Compressed Sparse Column (DCSC) format to store matrices. However, CombBLAS does not currently support GPU computation.
In contrast, GALATIC (Lett_2021) provides local SpGEMM on GPU and support for semiring abstraction but does not implement distributed memory parallelism. By extending the existing AC SpGEMM library (acspgemm), GALATIC achieves a significant speed-up compared to the local SpGEMM algorithms of CombBLAS. GALATIC uses a four-stage process for its operation. The process begins by dividing the non-zeros of $\mathbf{A}$ evenly among the blocks of a GPU. Then, the products are computed, expanded, sorted, and combined on the individual GPU threads. In the third step, parts of the product matrix are combined. The rows are then merged across the parts before the matrix is returned. This division of work makes for an extremely efficient SpGEMM, resulting in an algorithm that is faster than cuSPARSE for highly sparse matrices (acspgemm).
@@ 3. Related Work
Buluç et al. (bulucc2012parallel) developed and compared several distributed SpGEMM algorithms, including the Sparse SUMMA algorithm used in this work. Azad et al. (azad2016exploiting) presented a communication avoiding distributed SpGEMM algorithm that uses a 3D process grid and has a lower communication cost compared to Sparse SUMMA. Hong et al. (hong2024sparsity) presented a sparsity-aware distributed SpGEMM algorithm that reduces the communication cost by communicating only the non-zeros involved in the computation. These works do not consider the use of GPUs or explore hybrid communication.
Brock et al. (brock2024rdma) present a number of RDMA-based algorithms for distributed SpGEMM on GPUs, but do not investigate arbitrary semirings or hybrid communication. A large body of existing work (dalton2015optimizing; nagasaka2017high; acspgemm) focuses on optimizing SpGEMM on single GPUs. Other works (park2023orchestrating; liu2015framework) present strategies for using CPUs and GPUs on the same node in lockstep to perform SpGEMM. However, these works do not target distributed or multi-GPU scenarios.
@@ 4.3. GALATIC SpGEMM
Once the preparation is complete, the multiplication routine is called. GALATIC is highly configurable and has parameters for threads, threads per nonzero, and other ratios. In general, we have found that tuning these parameters does not have a significant impact on performance, but above a certain size and nonzero density, crashes can occur if these parameters are not set. The MaxChunksGeneralizedMerge setting has been adjusted to handle denser matrices by doubling the allowed matrix density for some shared memory segments. The configuration can be customized to the user’s need, e.g., performance or other behavior. If both input matrices have at least one nonzero, the multiplication $\mathbf{B}$ $\mathbf{A} T$ is performed. The resulting matrix is copied to the CPU, the GPU matrix is freed, and the CPU matrix is returned.
@@ 5. Results and Discussion
The experiments were performed on the GPU nodes of the Perlmutter Supercomputer at NERSC (nersc) using Cray-MPICH and CUDA Toolkit 12.4. Each GPU node is equipped with an AMD EPYC 7763 CPU, four NVIDIA A100 GPUs and four Cassini-1 Network Interface Cards (NICs). The GPUs on the same node are fully connected via 4 NVLink 3.0. The different nodes are connected via a 3-hop Dragonfly topology with an HPE Slingshot 11 interconnect.
The GPUs and CPUs are warmed up with an algorithm run that is not included in the statistics. I/O is excluded from the specified runtimes. The values for the first run can be found in Table 1. In GALATIC we used the default settings, except for MaxChunksGeneralizedMerge, which was set to 512 instead of 256 to handle larger matrices. The tests were performed with different matrices from the SuiteSparse Matrix Collection (davis2019algorithm), where each operation was a multiplication of a matrix by itself over the semiring of float types. Relevant statistics can be found in Table 2.
@@ 5.1. Distributed GPU-Enabled SpGEMM
For each of the four matrices, we performed three experiments, each of which was repeated four times across four different process grid sizes. For external comparison, we used the PETSc library(petsc-user-ref). PETSc offers 1D SpGEMM with row-wise partitioning and GPU acceleration. For comparison, we tested each matrix using CombBLAS CPU, CombBLAS GPU (this work), and PETSc GPU and performed $\mathbf{A}^{2}$ for each matrix. The tests ran on four nodes, with one GPU per MPI process. Our GPU routine used only CUDA-aware MPI broadcasts during these tests. The PETSc results in Figure 6 and the one process experiment in Figure 5 were excluded due to out of memory issues with large matrices. Our GPU-enabled CombBLAS version showed a significant speedup, although the magnitude varied depending on the number of processes. The highest speedup was $2\times$ on the Long_Coup_dt0 matrix. Overall, PETSc was faster with smaller matrices, but is slower or unable to run on larger matrices.
@@ 6. Conclusions
[…]